OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±7)

OpenStudy (jdoe0001):

the vertices are at (0, 2) and (0, -2) half-way between the vertices, is the center so your center is at (0, 0), or the origin from there goes UP 2 units and DOWN 2 units, so a = 2, and \(\bf a^2 = 4\) the foci are at (0, 7) and (0, -7) the distance from the center, (0, 0) to either of the foci is "c" $$ \bf c = \sqrt{a^2+b^2}\\ c^2 = a^2+b^2\\ (7)^2 = (2)^2+b^2\\ 49-4 = b^2 \implies b^2 = 45 $$

OpenStudy (jdoe0001):

the hyperbola is going up and down, that means the "y" rational will be the POSITIVE one